# Viscosity as the product of its low-concentration value times a thermodynamic function

### v1

@@ Abstract
The behavior of viscosity, $\eta$, as a function of concentration in dense fluids remains an unsolved problem, as is the case with other transport coefficients. Boltzmann’s theory and the Chapman-Enskog method predict the value of the viscosity at low concentrations, $\eta_{0}$. Here, the hypothesis $\eta=\phi\,\eta_{0}$ is proposed, where $\phi$ is a thermodynamic function representing the effects of interactions as concentration increases. The method proposed to verify this hypothesis involves coupling the system with a solvent represented by a Langevin thermostat, characterized by a damping time $t_{d}$. Molecular dynamics simulations show that different values of noise intensity modify $\eta$ and $\eta_{0}$, but do not affect $\phi$. This result supports the assumption that $\phi$ is a thermodynamic function, since the thermodynamic state remains unaltered by the presence of damping and noise. Simulations were conducted for particles that interact via a pseudo-hard sphere or a Lennard-Jones potential.
@@ I Introduction
The study and understanding of transport phenomena are highly relevant to general research and to many areas of industry. Transport features of fluids can be categorized into three main types. Diffusivity relates to the displacement of particles or molecules throughout the fluid, typically composed of particles of the same type. Thermal conductivity accounts for the exchange of kinetic energy between the constituents of the fluid that arises from collisions. Finally, viscosity quantifies the friction that occurs between two adjacent layers of a fluid where the molecules move on average with different velocities. More precisely, the viscosity coefficient indicates the momentum transfer rate between layers. For dilute gases, closed expressions can be derived using the Chapman-Enskog theory Chapman and Cowling (1970), which is based on Boltzmann’s equation. An analytical description covering the entire density range has proven to be a formidable challenge. There have been several attempts to address this issue, particularly for viscosity and other transport coefficients. Enskog extended his theory to dense fluids of hard spheres Enskog (1922), where the viscosity depends on the radial distribution function $g(\sigma)$ and the second virial coefficient $b$; $\sigma$ represents the diameter of the spheres. However, this expression is only valid for small to moderate densities because one of the assumptions in the formulation is the molecular chaos approximation. Koo and Hess derived an expression for dense hard sphere fluids using the Kirkwood-Smoluchowski equation Koo and Hess (1987), which agrees with simulation results but fails at low densities. Other authors have attempted to adapt the theory of Enskog to make it suitable for real fluids Hanley et al. (1972); the result is known as Modified Enskog Theory (MET). It assumes that a real fluid generally exhibits a behavior similar to that of a hard sphere fluid, specially for small or moderate concentrations. The values of $g(\sigma)$ and $b$ are obtained from existing equations of state for real fluids. In this context, the viscosity of hard spheres, $\eta_{HS}$, is related to the Enksog viscosity, $\eta_{E}$, by the expression $\eta_{HS}=F(\rho^{*})\eta_{E}$, where $\rho^{*}$ is the reduced density and $F(\rho^{*})$ is a correction function. Additionally, the HS viscosity can be expressed as $\eta_{HS}=G(\rho^{*})\eta_{0}$, where $\eta_{0}$ is the viscosity at low concentration and $G(\rho^{*})=F(\rho^{*})\eta_{E}/\eta_{0}$. The dense gas and liquid regions have been analyzed by Dymond Dymond (1974a, b, 1985); Dymond and Assael (1996). Different expressions for the correction function have been proposed by Silva Silva et al. (2003) and by Heyes and Sigurgeirsson Heyes and Sigurgeirsson (2004). The effective Hard-Sphere Diameter method represents a real fluid as a purely repulsive hard sphere system and then considers the attractive forces as a perturbation. For instance, the Weeks-Chandler-Andersen perturbation theory (WCA) Weeks et al. (1971) splits the Lennard-Jones potential into a reference term that contains only the repulsive forces, and a perturbation term that accounts for the attractive forces that manifest in a real fluid. Silva et al. provided an approximate expression for the viscosity of a fluid with WCA interactions Silva et al. (2003)
@@
A different method that yields good results involves excess entropy scaling laws. Rosenfeld Rosenfeld (1977a, b) derived a relation between the reduced viscosity $\eta^{*}$ and the excess entropy $S^{\text{ex}}$: $\eta^{*}=0.2\,\text{exp}(-0.8\,S^{\text{ex}}/Nk_{B})$, where $N$ is the number of particles and $k_{B}$ is the Boltzmann constant (more details on how to calculate reduced dimensionless quantities are provided below). This relation has proven to be quasi-universal, showing good agreement with simulations for various types of pair potentials, particularly in the moderate and large density regimes.
@@
As we see, a significant effort has been put to provide insight into the transport process theory. So far however, finding general expressions for the transport coefficients valid for the whole density range has proven to be an arduous task. In this manuscript we aim to better understand the role that interactions play in describing viscosity for dense systems. In Ref. Marchioni et al. (2023) it was suggested that the self-diffusion coefficient $D$ may be written as $D=\varphi D_{0}$, where $D_{0}$ is the self-diffusion coefficient at low densities (which can be calculated via the kinetic theory), and $\varphi$ is a function that only depends on the thermodynamic state of the system. This implies that microscopic details such as the mass or size of the particles are not relevant for $\varphi$; instead, $\varphi$ encapsulates the macroscopic information of the interactions. Here, we seek to verify whether an equivalent relation for the shear viscosity also holds:
@@
where $\eta_{0}$ is the low-density viscosity and $\phi$ is an unknown thermodynamic function.
The procedure is analogous to that described in Marchioni et al. (2023). We assume a system of interacting particles subjected to a Langevin thermostat. That is, the particles are immersed in a solvent composed of smaller ones, which apply both a stochastic force and a damping force. The idea is that the presence of noise should not have an effect on function $\varphi$, since we assume that it is a thermodynamic function, and noise does not modify the thermodynamic state. Conversely, noise intensity is expected to affect the low concentration viscosity $\eta_{0}$, as it depends on the microscopic details of the model.
We performed molecular dynamics simulations using LAMMPS software Plimpton (1995), in which the Stoermer-Verlet time integration algorithm for constant particle number, volume, and energy (NVE ensemble) is used.
@@
II The hypotheses and the method
In Ref. Marchioni et al. (2023), it was proposed that the self-diffusion coefficient obeys a general equation of the form $D=\varphi D_{0}$; the proposal was motivated by earlier work on transition probabilities of tagged particles between two neighboring small cells Di Muro and Hoyuelos (2021). It was demonstrated Marchioni et al. (2023) that numerical simulations are consistent with the hypothesis that $\varphi$ is a thermodynamic function, independent of microscopic details. Since this is an important guide for the development of a general transport theory for dense fluids, a natural step is to verify whether s similar equation holds for other transport coefficients. Here, we propose that viscosity conforms to a similar form, as given by Eq. (1), with $\phi$ a thermodynamic function.
@@ II The hypotheses and the method
where $t_{d}$ is the damping time, $m$ is the particle mass and $\mathbf{v}_{i}$ is the particle velocity; the term $-m\mathbf{v}_{i}/t_{d}$ represents friction and $\bm{\xi}_{i}$ is a stochastic force with zero mean and delta time correlated. If the temperature and the damping time are known, the noise intensity is determined by the fluctuation-dissipation theorem. The coupling to the thermostat is effectively represented by the damping time $t_{d}$.
The introduction of the thermostat modifies the transport properties; both $\eta$ and $\eta_{0}$ depend on $t_{d}$. However, the thermodynamic state remains unchanged. The equation of state is typically represented by the compressibility factor $Z=P/(k_{B}T\rho)$, where $P$ is the pressure, $k_{B}$ is Boltzmann constant, $T$ is the temperature and $\rho=N/V$ is the particle density. For given values of $T$ and $\rho$, the compressibility factor is determined by the pressure $P$. Using the virial equation (see, for example, Sec. 3.7 in Pathria and Beale (2011)), it can be shown that the pressure does not change when the thermostat is introduced (see Sec. IV in Marchioni et al. (2023) for more details); therefore, the thermodynamic state is independent of $t_{d}$. Since $\phi$ is assumed to be a thermodynamic function, it does not depend on $t_{d}$. Consequently, according to Eq. (1), $\eta/\eta_{0}$ is independent of $t_{d}$, unlike $\eta$ and $\eta_{0}$.
@@ III Viscosity at small concentration with Langevin thermostat
The small concentration viscosity can be analytically obtained using Boltzmann’s theory for the under-damped case (large $t_{d}$) and the Langevin description of Brownian motion for the over-damped case (small $t_{d}$). As discussed in the following sections, both pseudo-hard sphere and Lennard-Jones (LJ) potentials have a characteristic particle diameter, $\sigma$, and a characteristic energy, $\epsilon$. It is convenient to use reduced (dimensionless) quantities, denoted with an asterisk, which are obtained by scaling with a function of $\sigma$, $\epsilon$ or $m$. The reduced damping time is $t_{d}^{*}=t_{d}\,\sigma^{-1}\sqrt{\epsilon/m}$. Thus, a large $t_{d}$ means $t_{d}^{*}\gg 1$ and a small $t_{d}$ corresponds to $t_{d}^{*}\ll 1$.
@@
where $\Omega_{22}$ is the so-called reduced collision integral.
The reduced temperature is given by $T^{*}=k_{B}T/\epsilon$ and the reduced viscosity is $\eta^{*}=\eta\sigma^{2}/\sqrt{m\epsilon}$. The reduced collision integral for hard spheres (HS) is equal to 1, while for the LJ potential, it has been numerically shown Fokin et al. (1999) that it can be represented by the following function of the reduced temperature:
@@
with $a_{0}=0.31081$, $a_{1}=-0.171211$, $a_{2}=-0.715805$, $a_{3}=2.48678$, $a_{4}=-1.78317$, $a_{5}=0.394405$. Then, we have
@@
On the other hand, if $t_{d}^{*}\ll 1$ the interaction with the solvent particles (the reservoir) becomes significantly more dominant than the interaction between the larger particles in the system. In this case, the dynamics of the system closely resembles that of a Brownian particle. Using Langevin theory and the Green Kubo formula, the following closed expression can be derived (see Apendix A),
$\eta_{L}=\frac{\rho\,k_{B}T\,t_{d}}{2},$ (6)
or, using reduced quantities,
$\eta_{L}^{*}=\frac{\rho^{*}\,T^{*}\,t_{d}^{*}}{2}.$ (7)
This result represents the viscosity of the system of particles, which should not be confused with the viscosity of the solvent, which is related to the damping time. In fact, the damping time and the system viscosity are inversely proportional to the viscosity of the solvent.
For intermediate values of $t_{d}^{*}$, we employ a similar approach to that presented in Ref. Marchioni et al. (2023) for the self-diffusion coefficient. Since viscosity measures the capacity for momentum transfer between different parts of a fluid, its inverse represents resistance to this transfer. Momentum transfer can be described using Newton’s law of viscosity. For example, in laminar flow in the $x$ direction, $\tau_{xy}=\eta\,\frac{\partial{u}}{\partial{y}}$, where $\tau_{xy}$ is the shear stress (the force per unit area in the $x$ direction, also interpreted as the momentum density current in the $y$ direction), and $u$ is the average velocity of the fluid in the $x$ direction. This can be analogized to Ohm’s law in an electrical circuit, that is, $I=R^{-1}\mathcal{V}$, where $I$ is the electric current, $\mathcal{V}$ is the voltage and $R$ is the electric resistance. In this analogy, we make the following correspondences: $I\rightarrow\tau$, $\mathcal{V}\rightarrow\frac{\partial{u}}{\partial{y}}$, and $R\rightarrow\eta^{-1}$. By assuming that momentum flux is analogous to the electric current in a series circuit and considering the equivalent resistance, we can derive an expression for the low concentration viscosity:
$\frac{1}{\eta_{0}^{*}}=\frac{1}{\eta_{B}^{*}}+\frac{1}{\eta_{L}^{*}},$ (8)
In Fig. 1 we compare Eq. (9) with numerical simulations performed at low concentration for pseudo hard-spheres ($\Omega^{\text{HS}}_{22}=1$); we observe that there is a good agreement between theory and simulations.
Figure 1: Reduced viscosity at small concentration for pseudo hard-spheres, $\eta_{0}^{*}$, against damping time, $t_{d}^{*}$, for reduced density $\rho^{*}=0.01$. Crosses are numerical results and the curve corresponds to Eq. (9).
We can see in Fig. 2 that Eq. (9) also agrees well with the results of the simulations for intermediate values of damping time and for the LJ potential, where $\Omega^{\text{LJ}}_{22}$ is given by (4).
a)Refer to caption b)Refer to caption
[…]
IV Pseudo Hard Spheres
The pairwise potential for hard-spheres (HS) is
where $\sigma$ is the diameter of the spheres and $r$ is the distance between the centers of the spheres. It is well known that such interactions cannot be successfully implemented in Molecular Dynamics simulations, as particles that come very close to each other would experience an infinite force. However, it is possible to approximately model a hard-sphere fluid using Molecular Dynamics. Jover et al Jover et al. (2012) demonstrated that the Mie potential can effectively reproduce the properties of a hard spheres fluid. The potential is given by
$u(r)=\left\{\begin{array}[]{cc}\lambda_{r}(\frac{\lambda_{r}}{\lambda_{a}})^{\lambda_{a}}\epsilon\left[(\frac{\sigma}{r})^{\lambda_{r}}-(\frac{\sigma}{r})^{\lambda_{a}}\right]+\epsilon&\ \ \ r<\sigma\frac{\lambda_{r}}{\lambda_{a}}\\[6.0pt] 0&\ \ \ r\geq\sigma\frac{\lambda_{r}}{\lambda_{a}}\end{array}\right.,$ (11)
We have examined how viscosity varies with damping time at low concentration; see Eq. (9) and Fig. 1. The results indicate that, as the coupling with the reservoir decreases (i.e. as $t_{d}$ increases), the viscosity rises. At higher density, Fig. 3 illustrates the reduced viscosity as a function of reduced density for three different values of damping time (with the noise-free case corresponding to the limit of infinite $t_{d}$). As anticipated, for a given density, viscosity increases with larger $t_{d}$.
Figure 3: Viscosity for pseudo hard-spheres, $\eta^{\ast}$, versus reduced density, $\rho^{*}$, for different values of damping time $t_{d}^{*}$. The red crosses are noise-free results, blue crosses correspond to $t_{d}^{*}=10$ and green crosses to $t_{d}^{*}=100$. The circles indicate the numerical results from Pieprzyk et al. Pieprzyk et al. (2019).
Figure 4: Viscosity divided by low-concentration viscosity for pseudo hard-spheres, $\eta/\eta_{0}$, versus reduced density, $\rho^{*}$. See Fig. 3 for a description of the symbols.
In Fig. 4 we present simulation results of $\eta/\eta_{0}=\phi$ as a function of concentration for various damping times. We find that $t_{d}$ does not influence the behavior of $\eta/\eta_{0}$. The values of $\eta/\eta_{0}$ for different damping times $t_{d}^{*}$ agree within numerical errors. For comparison, we also include numerical results from Pieprzyk et al. Pieprzyk et al. (2019) for hard spheres (without noise).
V Lennard-Jones potential
where $\epsilon$ represents the depth of the potential well and $\sigma$ is the distance at which the potential is zero. We aim to investigate whether $\eta/\eta_{0}$ also behaves as a state function for a Lennard Jones fluid.
a)Refer to caption b)Refer to caption
Figure 5: Reduced viscosity $\eta^{\ast}$ for the Lennard-Jones potential as a function of the reduced density $\rho^{\ast}$, for temperatures (a) $T=1.5$ and b) $T=4$. The red crosses are noise-free results, blue crosses correspond to $t_{d}^{*}=10$ and green crosses to $t_{d}^{*}=100$.
Next, in Fig. 6 we divide the viscosity, $\eta$, by the viscosity at low concentration, $\eta_{0}$, as given by Eq. (9); that is, $\eta/\eta_{0}=\phi$. The resulting curves for different damping times collapse, confirming the hypotheses that $\phi$ depends solely on the thermodynamic state of the system.
a)Refer to caption b)Refer to caption
Figure 6: Viscosity divided by viscosity at small concentration, $\eta/\eta_{0}$, as a function of the reduced density $\rho^{\ast}$ for the Lennard-Jones potential, with temperatures a) $T=1.5$ and b) $T=4$. See Fig. 5 for a description of the symbols.
VI Conclusions
In this work, we investigated the hypothesis that the shear viscosity of dense fluids can be expressed as $\eta=\phi\,\eta_{0}$, where $\eta_{0}$ is the low-density viscosity and $\phi$ is a thermodynamic function.
We developed a method to test this hypothesis by coupling the fluid to a Langevin thermostat characterized by a damping time $t_{d}$. This allowed us to modify $\eta$ and $\eta_{0}$ without altering the thermodynamic state of the system. An analytical expression for $\eta_{0}$ as a function of $t_{d}$ was derived by combining Boltzmann and Langevin theories. This expression showed good agreement with numerical simulations for both pseudo hard spheres and Lennard-Jones fluids. Molecular dynamics simulations were performed for pseudo hard spheres and Lennard-Jones fluids over a range of densities and damping times. The results demonstrated that while $\eta$ depends on $t_{d}$, the ratio $\eta/\eta_{0}$ is independent of $t_{d}$ within numerical uncertainty. This invariance of $\eta/\eta_{0}$ with respect to $t_{d}$ supports our hypothesis that $\phi$ is indeed a thermodynamic function, as it depends only on the thermodynamic state of the system and not on microscopic details. The behavior was consistent for both pseudo hard spheres and Lennard-Jones potentials at different temperatures, suggesting the relation may be general for simple fluids.
Acknowledgments
Appendix A
@@ Acknowledgments
We consider a low-density system of Brownian particles, where interactions between particles are negligible. We aim to calculate the viscosity of this system. As mentioned in the text, it should not be confused with the viscosity of the medium in which the particles move, which is related to the friction coefficient $\gamma=1/t_{d}$, where $t_{d}$ is the damping time.
@@
Using this expression in Eq. A we obtain

### v2

# Viscosity as the product of its ideal low-concentration value times a thermodynamic function

@@ Abstract
The behavior of viscosity, $\eta$, as a function of concentration in dense fluids remains an unsolved problem, as is the case with other transport coefficients. Boltzmann’s theory and the Chapman-Enskog method predict the value of the viscosity at low concentrations, $\eta_{0}$. Here, the hypothesis $\eta=\phi\,\eta_{0}$ is proposed, where $\phi$ is a function of the thermodynamic state that represents the effects of interactions as concentration increases. We consider that $\eta_{0}$ is the viscosity in an ideal hypothetical system, where the condition of small interactions applies for the whole density range ($\phi\to 1$ for low concentration). The method proposed to verify this hypothesis involves coupling the system with a solvent represented by a Langevin thermostat, characterized by a damping time $t_{d}$. Molecular dynamics simulations show that different values of noise intensity modify $\eta$ and $\eta_{0}$, but do not affect $\phi$. This result supports the assumption that $\phi$ is a state function, since the thermodynamic state remains unaltered by the presence of damping and noise. Simulations were conducted for particles that interact via a pseudo-hard sphere or a Lennard-Jones potential.
@@ I Introduction
The study and understanding of transport phenomena are highly relevant to general research and to many areas of industry. Transport features of fluids can be categorized into three main types. Diffusivity relates to the displacement of particles or molecules throughout the fluid, typically composed of particles of the same type. Thermal conductivity accounts for the exchange of kinetic energy between the constituents of the fluid that arises from collisions. Finally, viscosity quantifies the friction that occurs between two adjacent layers of a fluid where the molecules move on average with different velocities. More precisely, the viscosity coefficient indicates the momentum transfer rate between layers. For dilute gases, closed expressions can be derived using the Chapman-Enskog theory Chapman and Cowling (1970), which is based on Boltzmann’s equation. An analytical description covering the entire density range has proven to be a formidable challenge. There have been several attempts to address this issue, particularly for viscosity and other transport coefficients. Enskog extended his theory to dense fluids of hard spheres Enskog (1922), where the viscosity depends on the radial distribution function $g(\sigma)$ and the second virial coefficient $b$; $\sigma$ represents the diameter of the spheres. A dense fluid is said to be a system of particles with a sufficiently high density that intermolecular interactions become significant, which makes the behavior of the system to deviate from the ideal gas assumptions. However, Enskog’s theory is only valid for small to moderate densities because one of the assumptions in the formulation is the molecular chaos approximation.
Koo and Hess derived an expression for dense hard sphere fluids using the Kirkwood-Smoluchowski equation Koo and Hess (1987), which agrees with simulation results but fails at low densities. Other authors have attempted to adapt the theory of Enskog to make it suitable for real fluids Hanley et al. (1972); the result is known as Modified Enskog Theory (MET). It assumes that a real fluid generally exhibits a behavior similar to that of a hard sphere fluid, specially for small or moderate concentrations. The values of $g(\sigma)$ and $b$ are obtained from existing equations of state for real fluids. In this context, the viscosity of hard spheres, $\eta_{HS}$, is related to the Enksog viscosity, $\eta_{E}$, by the expression $\eta_{HS}=F(\rho^{*})\eta_{E}$, where $\rho^{*}$ is the reduced density and $F(\rho^{*})$ is a correction function. Additionally, the HS viscosity can be expressed as $\eta_{HS}=G(\rho^{*})\eta_{0}$, where $\eta_{0}$ is the viscosity at low concentration and $G(\rho^{*})=F(\rho^{*})\eta_{E}/\eta_{0}$. The dense gas and liquid regions have been analyzed by Dymond Dymond (1974a, b, 1985); Dymond and Assael (1996). Different expressions for the correction function have been proposed by Silva Silva et al. (2003) and by Heyes and Sigurgeirsson Heyes and Sigurgeirsson (2004). A different proposal to improve Enskog’s theory proposed by Beijeren and Ernst is known as Revised Enskog theory (RET)van Beijeren and Ernst (1973), in which a spatially correlated radial distribution function is used, so as to capture the properties of nonuniform systems. This approach has been used to obtain the transport coefficients for multicomponent mixtures de Haro and Cohen (1984) and hard sphere crystals Kirkpatrick et al. (1990). Recently, Pieprzyk et al. have evaluated the transport coefficients of a hard sphere fluid and crystal by intensive Molecular Dynamics simulations Pieprzyk et al. (2024); they found that there is a good agreement with the results predicted by the RET.
The effective Hard-Sphere Diameter method represents a real fluid as a purely repulsive hard sphere system and then considers the attractive forces as a perturbation. For instance, the Weeks-Chandler-Andersen perturbation theory (WCA) Weeks et al. (1971) splits the Lennard-Jones potential into a reference term that contains only the repulsive forces, and a perturbation term that accounts for the attractive forces that manifest in a real fluid. Silva et al. provided an approximate expression for the viscosity of a fluid with WCA interactions Silva et al. (2003)
@@
A different method that yields good results involves excess entropy scaling laws. Rosenfeld Rosenfeld (1977a, b); Dyre (2018) derived a relation between a dimensionless viscosity $\tilde{\eta}$ and the excess entropy $S^{\text{ex}}$: $\tilde{\eta}=0.2\,\text{exp}(-0.8\,S^{\text{ex}}/Nk_{B})$, where $N$ is the number of particles and $k_{B}$ is the Boltzmann constant (more details on how to calculate reduced dimensionless quantities are provided below). This relation has proven to be quasi-universal, showing good agreement with simulations for various types of pair potentials, particularly in the moderate and large density regimes.
@@
As we see, a significant effort has been put to provide insight into the transport process theory. So far however, finding general expressions for the transport coefficients valid for the whole density range has proven to be an arduous task. In this manuscript we aim to better understand the role that interactions play in describing viscosity for dense systems. In Ref. Marchioni et al. (2023) it was suggested that the self-diffusion coefficient $D$ may be written as $D=\varphi D_{0}$, where $D_{0}$ is the self-diffusion coefficient at low densities (which can be calculated via the kinetic theory), and $\varphi$ is a function that depends only on the thermodynamic state of the system. This implies that microscopic parameters, such as the mass or size of the particles, do not explicity appear in general expressions of $\varphi$; instead, $\varphi$ encapsulates the macroscopic information of the interactions. Here, we seek to verify whether an equivalent relation for the shear viscosity also holds:
@@
where $\eta_{0}$ is the low-density viscosity and $\phi$ is an unknown function of the thermodynamic state.
The procedure is analogous to that described in Marchioni et al. (2023). We assume a system of interacting particles subjected to a Langevin thermostat. That is, the particles are immersed in a solvent composed of smaller ones, which apply both a stochastic force and a damping force. The idea is that the presence of noise should not have an effect on function $\phi$, since we assume that it is a thermodynamic function, and noise does not modify the thermodynamic state. Conversely, noise intensity is expected to affect the low concentration viscosity $\eta_{0}$, as it depends on the microscopic details of the model.
Cohen and de Schepper Cohen and de Schepper (1991), and also Heyes and Sigurgeirsson Heyes and Sigurgeirsson (2004), analyzed an analogy between a dense hard-sphere fluid and a colloidal suspension that uses as a key ingredient a relationship between the shear viscosity and self-diffusion. In a simulation, the solvent viscosity of the colloidal suspension can be progressively decreased giving place to a transition from Langevin to Newtonian dynamics Heyes and Sigurgeirsson (2004). An equivalent procedure is proposed here where, instead of decreasing the viscosity of the solvent, we consider a progressive increase of the damping time (Stokes formula indicates that the solvent viscosity is inversely proportional to the damping time). Nevertheless, there are important differences with our model. We neglect hydrodynamic forces (usually relevant in colloidal suspensions) and we do not assume the existence of a formal relationship between the shear viscosity and self-diffusion. The model that we use is a particle system (with hard sphere or Lennard-Jones interactions) with a Langevin thermostat; the feature that is interesting for us is that this model has an equation of state independent of the noise intensity, meaning that the thermodynamic state is not modified by the Langevin thermostat. The thermostat modifies the dynamics and the microscopic features, but not the thermodynamic state and, according to our hypothesis, not the function $\phi$.
We performed molecular dynamics simulations using LAMMPS software Plimpton (1995), in which the velocity form of the Stoermer-Verlet time integration algorithm for constant particle number, volume, and energy (NVE ensemble) is used (more details in appendix A).
@@
II The hypothesis and the method
In Ref. Marchioni et al. (2023), it was proposed that the self-diffusion coefficient obeys a general equation of the form $D=\varphi D_{0}$; the proposal was motivated by earlier work on transition probabilities of tagged particles between two neighboring small cells Di Muro and Hoyuelos (2021). It was demonstrated Marchioni et al. (2023) that numerical simulations are consistent with the hypothesis that $\varphi$ is a state function, independent of microscopic details. Since this is an important guide for the development of a general transport theory for dense fluids, a natural step is to verify whether s similar equation holds for other transport coefficients. Here, we propose that viscosity conforms to a similar form, as given by Eq. (1), with $\phi$ a state function.
@@ II The hypothesis and the method
where $t_{d}$ is the damping time, $m$ is the particle mass and $\mathbf{v}_{i}$ is the particle velocity; the term $-m\mathbf{v}_{i}/t_{d}$ represents friction and $\bm{\xi}_{i}$ is a stochastic force with zero mean ($\langle\xi_{i}(t)\rangle=0$) and delta time correlated ($\langle\xi_{i}(t)\xi_{i}(t^{\prime})\rangle=\frac{2mk_{B}T}{t_{d}}\delta(t-t^{\prime})$). If the temperature and the damping time are known, the noise intensity is determined by the fluctuation-dissipation theorem. The coupling to the thermostat is effectively represented by the damping time $t_{d}$.
The introduction of the thermostat modifies the transport properties; both $\eta$ and $\eta_{0}$ depend on $t_{d}$. However, the thermodynamic state remains unchanged. The equation of state is typically represented by the compressibility factor $Z=P/(k_{B}T\rho)$, where $P$ is the pressure, $k_{B}$ is Boltzmann constant, $T$ is the temperature and $\rho=N/V$ is the particle density. For given values of $T$ and $\rho$, the compressibility factor is determined by the pressure $P$. Using the virial equation (see, for example, Sec. 3.7 in Pathria and Beale (2011)), it can be shown that the pressure does not change when the thermostat is introduced (see Sec. IV in Marchioni et al. (2023) for more details); therefore, the thermodynamic state is independent of $t_{d}$. Since $\phi$ is assumed to be a thermodynamic function, it does not depend on $t_{d}$. To illustrate this point we show in Fig. 1 numerical results of the potential energy at equilibrium for different values of $t_{d}$ and different types of pair-wise interactions. We can observe that the noise intensity does not have an effect on the potential energy at equilibrium. Consequently, according to Eq. (1), $\eta/\eta_{0}$ is independent of $t_{d}$, unlike $\eta$ and $\eta_{0}$.
@@
a)Refer to caption b)Refer to caption
Figure 1: Dimensionless potential energy per particle $\mathcal{V}$ as a function of the reduced density $\rho^{*}$ (see Table 1 for a description of reduced quantities) for different values of reduced damping time $t_{d}^{*}$, of a system of $N$ particles interacting via pseudo-hard sphere (a) and Lennard Jones (LJ) potential (b). For the LJ potential, results for two values of the reduced temperature, $T^{*}=1.5$ and $T^{*}=4$, are shown. A number of particles $N=32000$ was used in the MD simulations. The statistical error of the data is not greater than $0.15\%$
@@ III Viscosity at small concentration with Langevin thermostat
The small concentration viscosity can be analytically obtained using Boltzmann’s theory for the under-damped case (large $t_{d}$) and the Langevin description of Brownian motion for the over-damped case (small $t_{d}$). As discussed in the following sections, both pseudo-hard sphere and Lennard-Jones (LJ) potentials have a characteristic particle diameter, $\sigma$, and a characteristic energy, $\epsilon$. It is convenient to use reduced (dimensionless) quantities, denoted with an asterisk, which are obtained by scaling with a function of $\sigma$, $\epsilon$ or $m$ (see Table 1). Note that a large damping time $t_{d}$ means $t_{d}^{*}\gg 1$ and a small $t_{d}$ corresponds to $t_{d}^{*}\ll 1$.
Quantities Real units Reduced quantities
Time $\tau$ $\tau^{\ast}=\tau\,\sigma^{-1}\sqrt{\epsilon/m}$
Temperature $T$ $T^{*}=k_{B}T/\epsilon$
Density $\rho$ $\rho^{*}=\rho\sigma^{3}$
Viscosity $\eta$ $\eta^{*}=\eta\sigma^{2}/\sqrt{m\epsilon}$
Table 1: Relations between real and reduced (dimensionless) quantities
III.1 Viscosity in the noise-free limit
@@
where $\Omega_{22}$ is the so-called reduced collision integral. The reduced collision integral for hard spheres (HS) is equal to 1, while for the LJ potential, it has been numerically shown Fokin et al. (1999) that it can be represented by the following function of the reduced temperature:
@@
with $a_{0}=0.31081$, $a_{1}=-0.171211$, $a_{2}=-0.715805$, $a_{3}=2.48678$, $a_{4}=-1.78317$, $a_{5}=0.394405$. Using the definitions of dimensionless quantities in table 1, we have
@@
In the context of the excess entropy scaling Rosenfeld (1977a, b); Dyre (2018), the dimensionless viscosity is obtained in a different way; it is given by $\tilde{\eta}=\eta\,\rho^{-2/3}(mk_{B}T)^{-1/2}$. The proposal is similar to ours in that a state function, the excess entropy in this case, is also used to represent a transport coefficient. But there are important differences. We propose that the ratio $\eta/\eta_{0}$ is a function of the thermodynamic state. Excess entropy scaling postulates that $\tilde{\eta}$ is an exponential of the excess entropy with two adjustable parameters, where the so-called macroscopically reduced units are used; this reduced units vary with the thermodynamic state; space, time or energy are scaled with combinations of density, temperature and particle mass. We do not apply any such scaling and do not propose a specific form for $\eta/\eta_{0}$.
III.2 Viscosity in a Brownian system
@@
If $t_{d}^{*}\ll 1$ the interaction with the solvent particles (the reservoir) becomes significantly more dominant than the interaction between the larger particles in the system. In this case, the dynamics of the system closely resembles that of a Brownian particle. We use the Langevin theory and the Green Kubo formula to derive an expression for the viscosity in this limit.
@@
We consider a low-density system of Brownian particles, where interactions between particles are negligible. We aim to calculate the viscosity of this system. It should not be confused with the viscosity of the medium in which the particles move, which is related to the friction coefficient $\gamma=1/t_{d}$.
@@
Using this expression in Eq. III.2 we obtain